directions: write an equation of the line that is parallel to the given line and passes through the given point.
problem: y=2x+2,(3,2)
i dont understand how to work these type of problems out and i was wondering if this could be shown to me in steps so that i can work out the rest of the problems on my homework assignment.
thank you

Answer by rodriguezh2(15)

Whenever you are doing these types of problems, First make sure your equation is in y = mx + b form. Which yours is: y = 2x + 2
Remember (m) represents the slope (which is 2 in your equation)
And (b) represents the y-intercept (which is also 2 in your equation)
Parallel lines have slopes that are equal to each other. Therefore we can say that the slope of the line we are looking for is also 2.
We know that the line we are looking for passes through the point (3, 2)
Remember that 3 represents the x value and 2 represents the y value.
We can use the following equation: (y - y1) = m(x - x1)
(y - 2) = 2(x - 3) Now we multiply through the right side of the equation
(y - 2) = 2x - 6 Now we add 2 to both sides of the equation
y = 2x - 4
The above is the equation of the line you are looking for!
Hope this helped!